OpenStudy (anonymous):

hi i really dont get these equations and i would like some help plz 12x-15=8x+5

OpenStudy (amoodarya):

variable on left side = numbers on right side then simplify them

OpenStudy (amoodarya):

12x-15=8x+5 12x-8x=15+5 can you complete this ?

OpenStudy (anonymous):

ill try and see what i come out with . thanks

OpenStudy (amoodarya):

i solve an example 3x+5=2x-6 3x-2x=-5-6 x=-11 now you try like this operation

OpenStudy (anonymous):

okay thanks got it

OpenStudy (amoodarya):

lets check the answer

OpenStudy (anonymous):

x=5

OpenStudy (amoodarya):

it's ok

OpenStudy (anonymous):

ok